The mean lifetime of a sample of 100 light tubes produced by a company is found to be 1570 hours with standard deviation of 80 hours. Test the hypothesis that the mean lifetime of the tubes produced by the company is 1600 hrs.


1
Expert's answer
2021-01-12T00:55:40-0500

n=100, "\bar X=1570, \sigma=80"

We test the hypotheses ;

"H_0:\mu=1600"

"H_1:\mu≠1600"

Since the sample size is greater than 30 and the population standard deviation is known we use the z test.

"Z=\frac{X-\mu} {\frac{\sigma} {\sqrt n}}"

="\frac{1600-1570}{\frac{80}{\sqrt{100}}}"

=3.75

We take a level of significance of 1%.

The corresponding critical value is;

"Z_{\alpha/2}=2.5758"

If the test statistic is greater than the critical value, we reject the null hypothesis in favor of the alternative hypothesis.

|3.75|>2.5758

We reject the null hypothesis hypothesis in favor of the alternative hypothesis.

We are 99% confident that the population mean lifetime of bulbs is not equal to 1600 hours.